# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v2

@@ Abstract
The foundation model enables fast problem-solving without learning from scratch, and such a desirable adaptation property benefits from its adopted cross-task generalization paradigms, e.g., pretraining, meta-training, and finetuning. Recent advances in these paradigms show the crucial role of challenging tasks’ prioritized sampling in enhancing adaptation robustness and even improving sampling efficiency. However, scoring task difficulties exhausts massive task queries and requires intensive evaluation and computations, e.g., policy evaluations in Markov decision processes (MDPs) or inference with large backbone models. This work underscores the criticality of both adaptation robustness and learning efficiency, especially in scenarios where tasks are risky to collect or costly to evaluate. To this end, we present Model Predictive Task Sampling (MPTS) to establish connections between the task space and adaptation risk landscape for robust active task sampling. Technically, MPTS characterizes the task episodic information with a generative model and predicts optimization outcome, i.e., task-specific adaptation risk values, from posterior inference. The resulting risk learner amortizes expensive annotation, evaluation, or computation operations in task robust adaptation. Extensive experimental results show that MPTS can be seamlessly integrated into zero-shot, few-shot, and many-shot learning paradigms, increases adaptation robustness, and retains learning efficiency without affording extra cost. The code is available at the project site https://github.com/thu-rllab/MPTS.
@@ 1 Introduction
Figure 1: Framework of Model Predictive Task Sampling in Any-Shot Learning. a. The task episodic learning might involve MDPs, large models, or annotators in task construction for optimization. The adaptive machine learner can be a large model or RL policy. The MPTS contains posterior inference, adaptation outcome simulation, and task subset selection for any-shot learning. [Snow: freeze models; Fire: update models] b. MPTS amortizes the process, e.g., task batch construction, environment interaction, or large models’ inference at a lower computational budget with the risk learner. c. The risk learner takes in the coupled identifier and the exact adaptation risk to train under streaming variational inference. d. The risk learner simulates the adaptation outcomes of $\hat{\mathcal{B}}$ candidate identifiers, computes acquisition scores, and selects $-\mathcal{B}$ ones for next iteration.
Enhancing machine learning models’ generalization across various scenarios is a long-standing pursuit in artificial general intelligence. The rise of generative artificial intelligence provides a promising solution to this problem, which facilitates the development of foundation models 1, 2, 3. Unlike traditional models, which are specific to learning tasks and cannot sufficiently meet the demands of time-sensitive responses, e.g., online intelligent recommendation or consultancy, the foundation model exhibits the advantages of fast deployment in a wide range of tasks at a lower computational cost.
One secret behind foundation models’ success lies in the rapid adaptation capability obtained from pretraining, meta-training, and finetuning in the task distribution. Take meta-learning as an example; such a paradigm shares a similar cognition in human’s developing skills 4, where we consolidate rich experience into knowledge and quickly convert the few-shot guidelines to a specific skill. Technically, adaptation learning paradigms episodically update the problem-solving prior and leverage the extracted knowledge to unseen but related scenarios in a zero-shot, few-shot, or many-shot manner 1, 5. Such a mechanism benefits from slow learning on massive tasks and avoids learning from scratch when faced with similar tasks at test time.
@@
Currently, there have been several studies in enhancing adaptation, e.g., the neural scaling law in foundation models 6, 7 implies that cross-task generalization scales with model complexity, task capacity, and computational power. Recognized as a few-shot learner, large language models (LLMs) mandate the collection of large datasets as tasks to train for generalization 8. In robotics, a popular schema for generalist policies is diversifying the decision-making environments, e.g., Markov decision processes (MDPs), for the robot to interact and execute policy optimization 9. As a consequence, task episodic learning is commonly used in learning to adapt over task distributions, where the task batch is randomly resampled, e.g., from a uniform random sampler, to optimize over iteration (see Fig. 1/2b). Meanwhile, these distributions are induced by the task identifiers; for example, in Fig. 2a, the physics parameters randomize MDPs in domain randomization (DR) 9 and meta reinforcement learning (Meta-RL) 10.
@@
Meanwhile, adaptation robustness is catching increasing attention in the field 11, 12, and distribution shifts 13 at the task level exist ubiquitously in the real world. Some trivial yet frequently adopted sampling strategies, e.g., random sampling from a uniform distribution over tasks, are prone to encounter catastrophic failure in adaptation to risk-sensitive scenarios. For example, traffic accidents are more informative and crucial than usual cases in training autonomous driving systems 14. This inspires some advances in challenging tasks’ prioritized sampling strategies 14, 15, 16, which necessitates evaluating task difficulties. For the sake of improving adaptation robustness, these approaches 17, 18, 16, 19, 20 score the task samples after evaluation and prioritize difficult ones to optimize (see evaluate-rank-filter in Table 1). However, exactly evaluating task difficulties from either losses or gradients brings extra computational and annotation overhead. In LLMs, the hallucination stands as a lasting challenge, and the post-training phase sometimes relies on prohibitively intensive manual evaluation and precise human feedback such as millions of interactions, which serves for effective alignment 21, 22. In DR or Meta-RL, the agent has to interact with numerous MDPs to collect episodes collection and compute returns. All of these raise concerns about learning efficiency during task robust adaptation, particularly when the machine learner employs a large backbone or annotations are expensive.
@@
Encouraged by the pressing real-world necessities, we delve into robust active task sampling, which can potentially cut off extra costs arising from task construction, manual annotation, or the computational intensive evaluation of foundation models in the task space. In the presence of any-shot adaptation, it is desirable to seek a task sampling strategy that requires fewer learning resources but retains more deployment benefits such as robustness.
@@
Our brain can simulate the outcome of some unencountered cases based on continually accumulated trial experience without exactly putting actions into practice. The inherent cognitive system is robust in handling nontrivial cases and energy efficient when developing diverse skills through an implicit selection mechanism. Hence, it is also reasonable to model the optimization process of adaptation learning and simultaneously adjust the task sampling strategy from the predicted outcome as feedback. The rationale is also reflected in neuroscience 4, 23, 24; the prefrontal cortex resembles a meta-reinforcement learning system, implying the adaptation feedback also interleaves with the task query process. As a result, this work investigates the design of task robust sampling strategies with the guidance of a predictive model. We posit key insights in methodology: (i) Some variables, i.e., adaptation risk, are predictable to score task difficulties throughout episodic learning, which is informative to steer task selection; (ii) Generative modeling task-specific adaptation risk enables capturing risk landscapes with quantified uncertainty, which aligns optimization processes with robustness principles.
@@
To this end, we develop a new framework oriented to task episodic learning, namely Model Predictive Task Sampling (MPTS) to enable robust active task sampling in risk minimization. As displayed in Fig. 1a, our strategy exploits the sequential risk information to construct a lightweight surrogate model referred to as the risk learner, providing adaptation risk foresight across the task space and establishing connections between the adaptive machine learner, the task sampler, and the optimizer. Specifically, we view actively selecting task batches as planning from past experience with the assistance of the risk learner in MPTS. Regarding scoring task difficulties, Fig. 1b reveals the risk learner’s role in amortizing the expensive evaluation process on an arbitrary set of tasks conditioned on an adaptive machine learner. Fig. 1c shows that the risk learner is in a variational autoencoder (VAE) structure 25 and engages in adaptation risk generation via posterior inference. The acquisition function in Fig. 1d comprises worst-case performance and the predictive uncertainty into selection criteria to rate the task identifier candidate for filtering purposes.
@@
MPTS also draws inspiration from active inference 26, which involves perception, action, and learning phases in a loop for uncertainty minimization about the planning environment. In a similar way, this work conceptually characterizes task episodic learning by a risk generative model in Fig. 2b-c, and the task selection in the batch can be treated as online planning to derive a robust machine learner. Technically, MPTS specifies or infers the identifiers from the task distribution (see some instantiations in Fig. 2a), employs the risk learner to forecast any-shot adaptation risk and runs acquisition functions in a larger identifier batch to pick up the task batch for updating the machine learner. For practical optimization of the risk learner, we apply the streaming variational inference (VI) toward the risk generative model and execute the approximate inference towards the task-specific adaptation risk. In primary, our proposed MPTS enjoys several benefits in robust optimization:
@@
    Learning Efficiency. As the lightweight risk learner amortizes expensive evaluation parts, MPTS can diminish computational overhead and encourage exploration in the task space;
@@
    Framework Versatility. Leveraging the episodic optimization history, MPTS can roughly score the task difficulties in adaptation and allow seamless integration into robust zero-shot, few-shot, and many-shot learning.
This work performs extensive experiments in diverse task episodic learning scenarios, including few-shot sinusoid regression, few-shot image classification with multimodal foundation models, Metra-RL, robotic DR, and prompt-tuning multimodal foundation models. Evidence indicates that incorporating MPTS into any-shot adaptation learning can significantly promote adaptation robustness in most scenarios. Importantly, the risk learner empowers MPTS to circumvent obtaining evaluation results from complicated optimization steps. Under certain scenarios, MPTS accelerates the learning process at the cost of fewer learning resources compared with existing state-of-the-art (SOTA) robust adaptation methods.
Figure 2: Fundamental Concepts: Task Identifiers, Episodic Learning and Probabilistic Graphical Models. a. The task distribution is uniform and defined over meaningful identifiers $\bm{\tau}$. For example, the amplitude and the phase $[a,b]$ specifies a sinusoid curve to complete with five observed data points. Robots like Half-Cheetahs are asked to accomplish different locomotion tasks with varying masses and velocities. Some multimodal pattern recognition tasks’ identifiers are implicit but can be approximately obtained from pretrained models like CLIP 1. b. The episodic learning in the task space resamples a batch of tasks from $p(\tau)$ to optimize in each iteration. c. Here, the generative model includes grey units as observed variables and white ones as unobservable. The solid directed lines describe the generative model 27. We use the dash-directed lines to indicate the recognition model and approximate inference within autoencoding variational Bayes 25.
@@ 2 Adaptation and Robustness
We denote a task sample by $\tau\sim p(\tau)$, with $\mathcal{T}$ the task space. The identifier, a real-value vector $\bm{\tau}$ illustrated in Fig. 2a, specifies the task $\tau$ in the distribution. Let the task-specific risk function be $\ell:\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}\times\Theta\mapsto\mathbb{R}$, which evaluates the $\bm{\theta}\in\bm{\Theta}$ parameterized machine learner’s adaptation performance on the task $\tau$. Take few-shot regression as an instance, $\mathcal{D}_{\tau}^{S}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=1}^{K}$ refers to the support dataset available for fast adaptation, e.g., finetuning to obtain the model $p_{\bm{\theta}}(\bm{y}|\mathcal{D}_{\tau}^{S},\bm{x})$; while $\mathcal{D}_{\tau}^{Q}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=K+1}^{K+N}$ denotes the query dataset to evaluate the after-adaptation machine learner’s performance, e.g., $\ell=-\frac{1}{N}\sum_{i=1}^{N}\ln p_{\bm{\theta}}(\bm{y}_{i}|\mathcal{D}_{\tau}^{S},\bm{x}_{i})$.
@@
The tuple set $\left\{\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$ involves the sampled task identifiers $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$, the query and support dataset $\{\mathcal{D}_{\tau_{t,i}}:=\mathcal{D}_{\tau_{t,i}}^{S}\cup\mathcal{D}_{\tau_{t,i}}^{Q}\}_{i=1}^{\mathcal{B}}$, and the evaluated adaptation risk $\{\ell_{t,i}\}_{i=1}^{\mathcal{B}}$ in the $t$-th iteration. For convenience, we express the simplified risk history as $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$, which depends on $\bm{\theta}_{t}$.
@@ 2.1 Adaptation Risk Function and Task Filtering
During training, we evaluate $\ell$ on the query dataset $\mathcal{D}_{\tau}^{Q}$ conditioned on the machine learner $\bm{\theta}$, i.e., $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$. In robotics, one typical example can be DR 9, where $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$ denotes the negative return of trajectories collected from the policy $\bm{\theta}$. This evaluation is commonly seen in zero-shot learning without support information.
@@
The form of $\ell$ is specific to meta-learning methods. For instance, MAML 28 implements a bi-level optimization framework. In this context, $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})$ is expressed as $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta}_{\text{meta}}-\alpha\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{S};\bm{\theta}))$, where $\bm{\theta}_{\text{meta}}$ denotes the meta initialization, and the inside-bracket gradient term tailored to the task $\tau$ corresponds to finetuning the model with $\alpha$ the learning rate.
@@
Recent works 29, 19, 30 uncover compelling evidence that filtering some tasks to train under several criteria, e.g., ‘easy,’ ‘difficult,’ or ‘learnable’ in evaluation can enhance machine learners’ adaptation robustness and even improve task efficiency compared with uniform sampling strategies without curation. The ‘difficult’ prioritization criteria aligns with minimizing conditional value-at-risk ($\text{CVaR}_{\alpha}$) 31, a metric in evaluating robustness 32, 16, 33. These methods operate under an Evaluate-Rank-Filter mode in Table 1 and induce the batch filtering ratio $\alpha_{\text{train}}=1-\frac{\mathcal{B}}{\hat{\mathcal{B}}}\in[0,1)$, which describes the relative fraction of sampled task discarded in a batch.
Table 1: A Summary of Typical Risk Minimization Approaches. The previously mentioned ERM 34, GDRM 35, 36, 37, DRM 38, 16, 33, 29, 30, and MPTS are included in comparison. In DRM, the filtering criteria is to discard proportional easy tasks in a batch.

### v3

@@ Abstract
The foundation model enables general-purpose problem-solving and enjoys desirable rapid adaptation due to its adopted cross-task generalization paradigms, e.g., pretraining, meta-training, and finetuning. Recent advances in these paradigms show the crucial role of challenging tasks’ prioritized sampling in enhancing adaptation robustness. However, ranking task difficulties exhausts massive task queries to evaluate, thus computation and annotation intensive, which is typically unaffordable in practice. This work underscores the criticality of both adaptation robustness and learning efficiency, especially in scenarios where tasks are risky or costly to evaluate, e.g., policy evaluations in Markov decision processes (MDPs) or inference with large models. To this end, we present Model Predictive Task Sampling (MPTS) to establish connections between the task space and adaptation risk landscape to form a theoretical guideline in robust active task sampling. MPTS characterizes the task episodic information with a generative model and directly predicts task-specific adaptation risk values from posterior inference. The developed risk learner can amortize expensive evaluation and provably approximately rank task difficulties in the pursuit of task robust adaptation. MPTS can be seamlessly integrated into zero-shot, few-shot, and many-shot learning paradigms. Extensive experimental results are conducted to exhibit the superiority of the proposed framework, remarkably increasing task adaptation robustness and retaining learning efficiency in contrast to existing state-of-the-art (SOTA) methods. The code is available at the project site https://github.com/thu-rllab/MPTS.
@@ 1 Introduction
Figure 1: Framework of Model Predictive Task Sampling in Any-Shot Learning. a. The task episodic learning might involve MDPs, large models, or annotators in task construction for optimization. The adaptive machine learner can be a large model or RL policy. MPTS contains posterior inference, adaptation outcome simulation, and task subset selection for any-shot learning. [Snow: freeze models; Fire: update models] b. MPTS amortizes the process, e.g., task batch construction, environment interaction, or large models’ inference at a lower computational budget with the risk learner. c. The risk learner takes in the coupled identifier and the exact adaptation risk to train under streaming variational inference. d. The risk learner simulates the adaptation outcomes of $\hat{\mathcal{B}}$ candidate identifiers, computes acquisition scores, and selects $-\mathcal{B}$ ones for next iteration.
Enhancing machine learning models’ generalization across various scenarios is a long-standing pursuit in artificial general intelligence. The rise of generative artificial intelligence provides a promising solution to this problem, which facilitates the development of foundation models 1, 2, 3. Unlike traditional models, which are specific to learning tasks and cannot sufficiently meet the demands of time-sensitive responses, e.g., online intelligent recommendation or consultancy, the foundation model exhibits the advantages of fast deployment in a wide range of tasks without learning from scratch.
One secret behind foundation models’ success lies in rapid adaptation acquired from pretraining, meta-training, and finetuning. Take meta-learning as an example; it optimizes the machine learner over the task distribution and shares a similar cognition in human’s developing skills 4, where we consolidate cumulated experience into knowledge and quickly convert the few-shot guidelines to a specific skill. Technically, adaptation learning paradigms update machine learners episodically and leverage extracted problem-solving prior to unseen but related scenarios in a zero-shot, few-shot, or many-shot manner 1, 5. Such a mechanism exploits slow learning on massive tasks to solve new tasks fast at test time.
@@
Currently, there have been several studies in enhancing adaptation, e.g., the neural scaling law in foundation models 6, 7 implies that cross-task generalization scales with model complexity, task capacity, and computational power. Recognized as a few-shot learner, large language models (LLMs) mandate the collection of large datasets as tasks to train for generalization 8. In robotics, a popular schema for generalist policies is diversifying the decision-making environments, e.g., Markov decision processes (MDPs), for the robot to interact and execute policy optimization 9. As a consequence, task episodic learning is commonly used in learning to adapt over task distributions, where the task batch is randomly resampled, e.g., from a uniform random sampler, to optimize over iteration (see Fig. 1/2b). Meanwhile, these distributions are induced by the task identifiers; e.g., in Fig. 2a, diverse physics parameters are used to create diverse MDPs in domain randomization (DR) 9 and meta reinforcement learning (Meta-RL) 10.
@@
Meanwhile, as distribution shifts 11 exist ubiquitously at the task level in the real world, adaptation robustness is catching increasing attention in the field 12, 13. Some trivial yet frequently adopted sampling strategies, e.g., random sampling from a uniform task distribution, are prone to encounter catastrophic failure in adaptation to risk-sensitive scenarios due to undersampling critical tasks. For example, traffic accidents are more informative than usual cases when training autonomous driving systems but occupy a small portion 14. This inspires some advances in challenging tasks’ prioritized sampling strategies 14, 15, 16, which necessitates task difficulty ranking during optimization. For the sake of improving adaptation robustness, these approaches 17, 18, 16, 19, 20 rank the task samples after evaluation and prioritize difficult ones to optimize (see evaluate-rank-filter in Table 1). Nevertheless, exactly evaluating tasks from either losses or gradients brings extra computational and annotation overhead. In LLMs, the hallucination stands as a lasting challenge, and the post-training phase sometimes relies on prohibitively intensive manual evaluation and precise human feedback such as millions of interactions, which serves for effective alignment 21, 22. In DR or Meta-RL, the agent has to interact with numerous MDPs to collect after-adaptation episodes and compute returns. All of these raise concerns about learning efficiency during optimization for robust adaptation, particularly when the machine learner employs a large backbone or interactions with environments are expensive.
@@
Encouraged by the pressing real-world necessities, we delve into robust active task sampling, which can potentially cut off extra costs arising from task construction, intensive annotations, or computations when evaluating the machine learner’s adaptation to tasks. In the presence of any-shot adaptation, it is desirable to seek a task sampling strategy that requires fewer learning resources but retains more deployment benefits such as robust pattern recognition with foundation models and risk-averse sequential decision-making.
@@
Our brain can simulate the outcome of some unencountered scenarios based on continually cumulated experience without exactly putting actions into practice. The inherent cognitive system is also robust to handle extreme cases and energy efficient when developing diverse skills through an implicit task selection mechanism. Hence, it prompts us to learn to model the optimization process of any-shot adaptation and simultaneously adjust the task sampling strategy from the predicted outcome as feedback. A similar rationale is reflected in neuroscience 4, 23, 24; the prefrontal cortex resembles a meta-reinforcement learning system, implying the adaptation feedback also interleaves with the task query process. As a result, this work investigates the design of task robust sampling strategies with the guidance of a risk predictive model. We posit key insights in methodology: (i) Some variables, i.e., adaptation risk, are predictable during episodic learning, which is informative to rank task difficulties and steer task selection; (ii) Generative modeling task-specific adaptation risk enables capturing risk landscapes with quantified uncertainty, which aligns optimization processes with certain robustness principles.
@@
To this end, we develop a new framework oriented to task episodic learning, namely Model Predictive Task Sampling (MPTS) to enable robust active task sampling in risk minimization. As displayed in Fig. 1a, our strategy exploits the historical risk information to construct a lightweight surrogate model referred to as the risk learner, foresighting adaptation risk across the task space for the task sampler and guiding the optimization of the adaptive machine learner. Specifically, we view actively selecting task batches as planning from past experience with the assistance of the risk learner in MPTS. Fig. 1b reveals the risk learner’s role in amortizing the expensive evaluation process on an arbitrary set of tasks conditioned on an adaptive machine learner, which further serves task difficulty scoring. The risk learner in Fig. 1c is in a variational autoencoder (VAE) structure 25 and engages in online adaptation risk generation via posterior inference. The acquisition function in Fig. 1d comprises worst-case performance and the predictive uncertainty into selection criteria to rate the task identifier candidate for filtering purposes.
@@
MPTS also draws inspiration from active inference 26, which involves perception, action, and learning phases in a loop for uncertainty minimization about the planning environment. Similarly, this work conceptually characterizes task episodic learning by a risk generative model in Fig. 2b-c, and the subset selection from the task batch can be treated as online planning to derive a robust machine learner. Technically, MPTS specifies or infers the identifiers from the task distribution (see some instantiations in Fig. 2a) to establish mappings inside the risk learner and adopts the streaming variational inference (VI) in training the model. In addition, running acquisition functions in a larger identifier batch to pick up the task subset for the next iteration balances exploration (uncertainty minimization) and exploitation (worse-case robustness) in the task space. In primary, our proposed MPTS enjoys several benefits in robust optimization:
@@
    Learning Efficiency. As the lightweight risk learner amortizes expensive evaluation parts, MPTS can diminish computational overhead, avoid unnecessary annotations, and encourage exploration in the task space;
@@
    Framework Versatility. Leveraging the optimization episodes, MPTS can work as a plug-play module to rank the task difficulties in optimization and allow seamless integration into robust zero-shot, few-shot, and many-shot learning.
This work performs extensive experiments in diverse task episodic learning scenarios, including few-shot regression, few-shot image classification with foundation models, Metra-RL, robotic DR, and prompt-tuning foundation models. Evidence indicates that incorporating MPTS into any-shot adaptation learning can significantly promote adaptation robustness in most scenarios. Compared with existing SOTA robust adaptation methods, MPTS raises no extra computation, memory usage, or interactions with environments and even accelerates the learning process under certain scenarios.
Figure 2: Fundamental Concepts: Task Identifiers, Episodic Learning and Probabilistic Graphical Models. a. The task distribution is uniform and defined over meaningful identifiers $\bm{\tau}$. For example, the amplitude and the phase $[a,b]$ specifies a sinusoid curve to complete with K-shot observed data points. Robots like Half-Cheetahs are trained to accomplish different locomotion tasks with varying masses and velocities. Some multimodal pattern recognition tasks’ identifiers are implicit but can be described from a reference model, e.g., text encoders in CLIP 1. b. The episodic learning in the task space resamples a batch of tasks from $p(\tau)$ to optimize in each iteration. c. Here, the generative model includes grey units as observed variables and white ones as unobservable. The solid directed lines describe the generative model 27. We use the dash-directed lines to indicate the recognition model and approximate inference within autoencoding variational Bayes 25.
@@ 2 Adaptation and Robustness
We denote a task sample by $\tau\sim p(\tau)$, with $\mathcal{T}$ the task space. The identifier, a real-value vector $\bm{\tau}$ illustrated in Fig. 2a, specifies the task $\tau$ in the distribution. Let the task-specific risk function be $\ell:\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}\times\Theta\mapsto\mathbb{R}$, which evaluates $\bm{\theta}$ parameterized machine learner’s adaptation performance on the task $\tau$. Take few-shot regression as an instance, $\mathcal{D}_{\tau}^{S}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=1}^{K}$ refers to the support dataset available for fast adaptation, e.g., finetuning to obtain the model $p_{\bm{\theta}}(\bm{y}|\mathcal{D}_{\tau}^{S},\bm{x})$; while $\mathcal{D}_{\tau}^{Q}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=K+1}^{K+N}$ denotes the query dataset to evaluate the after-adaptation machine learner’s performance, e.g., $\ell=-\frac{1}{N}\sum_{i=1}^{N}\ln p_{\bm{\theta}}(\bm{y}_{i}|\mathcal{D}_{\tau}^{S},\bm{x}_{i})$.
@@
The tuple set $\left\{\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$ involves the sampled task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$, the query and support dataset $\{\mathcal{D}_{\tau_{t,i}}:=\mathcal{D}_{\tau_{t,i}}^{S}\cup\mathcal{D}_{\tau_{t,i}}^{Q}\}_{i=1}^{\mathcal{B}}$, and the evaluated adaptation risk $\{\ell_{t,i}\}_{i=1}^{\mathcal{B}}$ in the $t$-th iteration. For convenience, we express the simplified risk history as $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$, which depends on $\bm{\theta}_{t}$.
@@ 2.1 Adaptation Risk Function and Task Filtering
During training, we evaluate $\ell$ on the query dataset $\mathcal{D}_{\tau}^{Q}$ conditioned on the machine learner $\bm{\theta}$, i.e., $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$. With robotic DR 9 as an example, $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$ denotes the negative return of trajectories collected under the policy $\bm{\theta}$. This evaluation is commonly seen in zero-shot learning without support information.
@@
The form of $\ell$ is specific to meta-learning methods. For instance, MAML 28 implements a bi-level optimization framework. In this context, $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})$ is written as $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta}_{\text{meta}}-\alpha\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{S};\bm{\theta}))$, where $\bm{\theta}_{\text{meta}}$ denotes the meta initialization, and the inside-bracket gradient term corresponds to finetuning the model tailored to $\tau$ with $\alpha$ the learning rate.
@@
Recent works 29, 19, 30 uncover compelling evidence that filtering some tasks to train under several criteria, e.g., ‘easy,’ ‘difficult,’ or ‘learnable’ in evaluation can enhance machine learners’ adaptation robustness and even improve task efficiency compared with uniform sampling strategies. The ‘difficult’ prioritization criteria aligns with minimizing conditional value-at-risk ($\text{CVaR}_{\alpha}$) 31, a metric in evaluating robustness 32, 16, 33. These methods operate under an Evaluate-Rank-Filter mode for the next model update in Table 1 and induce the batch filtering ratio $\hat{\alpha}=1-\frac{\mathcal{B}}{\hat{\mathcal{B}}}\in[0,1)$, which describes the fraction of discarded tasks in a sample batch.
Table 1: A Summary of Typical Risk Minimization Approaches. The previously mentioned ERM 34, GDRM 35, 36, 37, DRM 16, 33, 29, 30, and MPTS are included in comparison. In DRM, the filtering criteria is to discard proportional easy tasks in a batch.